# Timelike entanglement entropy with gravitational anomalies

### v1

@@ 4.4 Holographic TEE at finite temperature and angular potential
where $\lambda$ parametrizes the geodesics, mirroring the similar role $t$ played in (4.24). The first equation represents the spacelike geodesics joining end points of interval and null infinities while second equation is the timelike geodesic which joins these two null infinities. Here $R_{0}$ is some constant and $R_{P}$ is the length of timelike interval $A$ in the Poincaré coordinates as
@@
For the anomalous contribution, we can compute it separately for contribution from spacelike and timelike geodesics, however doing it is very complicated in rotating BTZ background. To compute it more simply, we use the fact that geodesic comprising of spacelike and timelike geodesic forms a single continuous curve homologous to the boundary interval and the integral $\int_{C}\tilde{n}\cdot\nabla n$ is insensitive to variation of $n$ along the curve and, only depends on the initial and final value of normal vector $n$. Since the initial and final points of geodesics lie on the spacelike geodesic, the boundary conditions on the normal vector $n$ can be written in terms of time coordinate used to describe the black hole similar to [4]
@@
Here $\lambda$ plays the role of $t$ as discussed earlier and takes value at initial point $\lambda_{i}=-\frac{R_{P}}{2}$ and $\lambda_{f}=\frac{R_{P}}{2}$ at the final point. Now using the coordinate transformation (4.40) to transform $(q,\tilde{q})$ into the BTZ coordinates followed by (4.47) and (4.11), we obtain the anomaly contribution to holographic TEE as follows
@@ 4.5 Holographic TEE at zero temperature and finite angular potential
In the extremal limit, the Hawking temperature $T_{H}$ and left moving temperature $T_{L}$ vanishes but there remains an effective temperature called the Frolov-Thorne temperature that represents the degeneracy of the ground state of an extremal BTZ black hole and given by [67]
@@
The metric in (4.51) can be mapped to the pure AdS3 spacetime in the Poincaré coordinates (4.41) by the following coordinate transformations [74]
@@
We observe that the imaginary part corresponding to left moving mode is absent similar to earlier results and terms containing $c_{R}$ resembles the TEE at an effective Frolov-Thorne temperature [67]. It also matches with the dual field theory result given in (3.8).
@@ 5 Summary and discussion
Subsequently we computed the timelike entanglement entropy for a pure timelike interval in CFT2 with gravitational anomalies dual to bulk topologically massive gravity (TMG) in asymptotically AdS3 geometries. The bulk three dimensional action for the TMG-AdS3 geometries involve a gravitational Chern-Simons term in addition to the usual Einstein-Hilbert term. In the locally AdS3 geometry we considered, the worldlines of massive spining particles are geodesics. Since the holographic TEE consists of spacelike and timelike geodesics, it involves the on-shell actions of massive spinning particles moving on these extremal worldlines (geodesics) in the dual bulk geometry. We derived from the Chern-Simons action the anomalous contribution specifically arising from timelike geodesic which represents the rotation of normal frames along it. Therefore, the holographic TEE involves the sum of length of the geodesics and the twist (rotation) of the normal frame along these geodesics. We further proposed a prescription for the boundary condition on normal vectors to compute the Chern-Simons contribution to the holographic TEE. Utilizing this prescription, we determined the holographic TEE in a CFT2 dual to pure AdS3 and also for the cases of rotating extremal and non-extremal BTZ black holes. Interestingly we observed that the holographic results in all these cases agrees precisely with the dual field theory results.
There are several interesting directions to follow from our work. In this paper, we obtained the holographic TEE using spacelike and timelike geodesics. Exploring the same using the RT surface in a complexified geometry as proposed in [31] can offer new insights about the anomaly contribution. Furthermore, exploring TEE in the Chern-Simons formulation of $(2+1)$-dimensional topologically massive gravity theories [75] and geometries that deviate from locally AdS3 such as warped AdS [76] would be an interesting areas of investigation. Computing TEE in higher dimensional theories with gravitational anomalies will also be interesting. Finally it would be also interesting to investigate TEE in the context of boundary Lifshitz field theory [77] along the lines of AdS/BCFT [39]. These investigations can provide further new insights into black hole interiors and bulk reconstruction. We leave these interesting issue for future investigations.

### v2

@@ 4.4 Holographic TEE at finite temperature and angular potential
where $\lambda$ parametrizes the geodesics, mirroring the similar role $t$ played in (4.25). The first equation represents the spacelike geodesics joining end points of interval and null infinities while the second equation is the timelike geodesic which joins these two null infinities. Here $R_{0}$ is some constant and $R_{P}$ is the length of timelike interval $A$ in the Poincaré coordinates as
@@
For the anomalous contribution, we can compute it separately for contribution from spacelike and timelike geodesics, however doing it is very complicated in rotating BTZ background. To compute it more simply, we use the fact that geodesic comprising of spacelike and timelike geodesic forms a single continuous curve homologous to the boundary interval and the integral $\int_{C}\tilde{n}\cdot\nabla n$ is insensitive to variation of $n$ along the curve and, only depends on the initial and final value of normal vector $n$. Since the initial and final points of geodesics lie on the spacelike geodesic, the boundary conditions on the normal vector $n$ can be written in terms of time coordinate used to describe the black hole similar to Castro:2014tta
@@
Here $\lambda$ plays the role of $t$ as discussed earlier and takes value at initial point $\lambda_{i}=-\frac{R_{P}}{2}$ and $\lambda_{f}=\frac{R_{P}}{2}$ at the final point. Now using the coordinate transformation (4.41) to transform (q,q~)($\tilde{q}$) into the BTZ coordinates followed by (4.48) and (4.12), we obtain the anomaly contribution to holographic TEE as follows
@@ 4.5 Holographic TEE at zero temperature and finite angular potential
In the extremal limit, the Hawking temperature $T_{H}$ and left moving temperature $T_{L}$ vanishes but there remains an effective temperature called the Frolov-Thorne temperature that represents the degeneracy of the ground state of an extremal BTZ black hole and given by PhysRevD.39.2125
@@
The metric in (4.52) can be mapped to the pure AdS3 spacetime in the Poincaré coordinates (4.42) by the following coordinate transformations Keski-Vakkuri:1998gmz
@@
We observe that the imaginary part corresponding to left moving mode is absent similar to earlier results and terms containing $c_{R}$ resembles the TEE at an effective Frolov-Thorne temperature PhysRevD.39.2125 . It also matches with the dual field theory result given in (3.8).
@@ 5 Summary and discussion
Subsequently we computed the timelike entanglement entropy for a pure timelike interval in CFT2 with gravitational anomalies dual to bulk topologically massive gravity (TMG) in asymptotically AdS3 geometries. The bulk three dimensional action for the TMG-AdS3 geometries involve a gravitational Chern-Simons term in addition to the usual Einstein-Hilbert term. In the locally AdS3 geometry we considered, the worldlines of massive spinning particles are geodesics. Since the holographic TEE consists of spacelike and timelike geodesics, it involves the on-shell actions of massive spinning particles moving on these extremal worldlines (geodesics) in the dual bulk geometry. We derived from the Chern-Simons action the anomalous contribution specifically arising from timelike geodesic which represents the rotation of normal frames along it. Therefore, the holographic TEE involves the sum of length of the geodesics and the twist (rotation) of the normal frame along these geodesics. We further proposed a prescription for the boundary condition on normal vectors to compute the Chern-Simons contribution to the holographic TEE. Utilizing this prescription, we determined the holographic TEE in a CFT2 dual to pure AdS3 and also for the cases of rotating extremal and non-extremal BTZ black holes. Interestingly we observed that the holographic results in all these cases agrees precisely with the dual field theory results.
There are several interesting directions to follow from our work. In this paper, we obtained the holographic TEE using spacelike and timelike geodesics. Exploring the same using the RT surface in a complexified geometry as proposed in Heller:2024whi can offer new insights about the anomaly contribution. Furthermore, exploring TEE in the Chern-Simons formulation of $(2+1)$-dimensional topologically massive gravity theories Ammon:2013hba and geometries that deviate from locally AdS3 such as warped AdS Detournay:2012pc would be an interesting areas of investigation. Determining the various inequalities satisfied by TEE similar to the entanglement entropy remains an important open problem. Computing TEE in higher dimensional theories with gravitational anomalies will also be interesting. Finally it would be also interesting to investigate TEE in the context of boundary Lifshitz field theory Chu:2024nwf along the lines of AdS/BCFT Chu:2023zah . These investigations can provide further new insights into black hole interiors and bulk reconstruction. We leave these interesting issue for future investigations.